OpenStudy (anonymous):

What is the solution for the equation 6x − 8 = 4x? x = ___. (Input whole number only). Numerical Answers Expected! Answer for Blank 1: Please help !!!!!!

OpenStudy (anonymous):

im pretty sure x=4 because you subtract four x and add 8 on both sides then divide by 2 to get four

OpenStudy (anonymous):

6x-8=4x +8 +8 6x=4x+8 -4x -4x 2x=8 divide both sides by two then you get 4

OpenStudy (anonymous):

okay thank youu .